OpenStudy (anonymous):

how do you expand ln (x^5/2y) ?

OpenStudy (anonymous):

x.x.x.x.x/2y you mean expanded form?

OpenStudy (anonymous):

yes

OpenStudy (shadowfiend):

You can use the rule of logarithms -- specifically, that division turns into subtraction: \[\ln{\frac{x^5}{2y}} = \ln x^5 - \ln 2y\]

OpenStudy (anonymous):

and that's it?

OpenStudy (shadowfiend):

I believe, but it depends on what `expand' means :) That's fairly typical, though.

OpenStudy (anonymous):

thanx a lot you rule!

OpenStudy (shadowfiend):

Glad to help! :)

OpenStudy (anonymous):

one more step. \[5ln(x)-ln(2y)\] \[2ln(x)-ln(y)-ln(2)\]

OpenStudy (anonymous):

typo in the last one. should be \[ln(\frac{x^5}{2y})=5ln(x)-ln(y)-ln(2)\]

OpenStudy (shadowfiend):

Damn, good call. satellite.